Sunday, July 1, 2012

Day 35, 7/1/2012 -- Couldn't think of a title!

Today started off not particularly well -- we had set the fridges to cycle overnight so we wouldn't be caught with a fridge that ran out in the middle of taking data. This all worked perfectly fine, except we noticed that one of the fridges started warming up immediately after we got in to the high bay in the morning. It was then that we realized we forgot a crucial step in the cycling process. Because the cryostat will, on average, live tilted 25 degrees off vertical during flight, the fridges inside are mounted so that they will be vertical when the cryostat is tilted 25 degrees. When the cryostat is NOT tilted, the fridge cycle...doesn't really work (due to the geometry of the particular fridges we use). And, well, we forgot to tilt the cryostat last night before we left.

So we ended up running the fridge cycle again this morning, at 9AM after tilting the cryostat. And the fridge cycle takes about 5.5 hours just to run...and then a few more hours before the focal planes are stable at their base temperatures, which is necessary before we can start working with the detectors. Which basically meant that for me and Jeff, our day was completely shot in terms of taking receiver calibration data. However, we did manage to make some parts to upgrade the Ebert-Fastie: we use a servo-controlled mechanical chopper wheel to modulate the signal coming out of the Ebert-Fastie. However, we noticed that the chop frequency wasn't always very stable, so Jeff had the idea to try coupling the chopper wheel to a stepper motor rather than a servomotor. This ended up being surprisingly easy, since Jeff had on hand some small stepper motors that had the same shaft size as the servomotor, and a spare stepper motor controller. All that needed to be done was to wire up the controller (Jeff's task) and make a new motor mount (my task). And, lo and behold, it all works! It looks like the chopper is MUCH more stable with a stepper motor driving it.

Elsewhere in the high bay, the ACS team continues to work on getting things ready to Bemco test. They have now reassembled the star cameras and disk pressure vessels and purged them with dry nitrogen before sealing them. And Andrei and KyleH did initial tests of wrapping the BTS with its aluminized mylar skin.

In the next few days, the other payloads headed to Antarctica will be showing up, so we had to move all of our miscellaneous stuff out of the west high bay and find a home for it in our high bay.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120701

No comments:

Post a Comment